OpenStudy (anonymous):

change the exential expression to an equielent expression invovling a log .. e^x=2 change the log equation to an equivalent equation invovling an exponent log 4^16=2

OpenStudy (anonymous):

e^x=2 log(e^x) = log(2) // you can pull out any exponent in a log. x*log(e) = log(2) //log(e) = 1 x = log(2)